# Constraints on the anomalous Higgs boson couplings in $Z\gamma\gamma$ production at muon collider

### v1

@@ III Analysis of Signal and Background Events
$\overline{c}_{\gamma}=1$ 193 0.95 188 1.09 184 1.16 149 1.91 94 13.8 75 300
$\widetilde{c}_{\gamma}=1$ 429 2.11 414 2.41 403 2.55 317 4.06 194 28.4 156 624
$\overline{c}_{HB}=1$ 41218 203 38484 224 37735 239 35343 453 28834 4221 23656 94624
$\widetilde{c}_{HB}=1$ 166 0.82 159 0.92 155 0.98 101 1.29 34 4.98 23 92.0
$\overline{c}_{HW}=1$ 106052 522 96585 561 95311 603 90336 1158 73853 10813 61391 245564
$\widetilde{c}_{HW}=1$ 221 1.09 212 1.23 207 1.31 162 2.08 97 14.2 83 332
Background
$Z\gamma\gamma$ 70 67 65 38 4.33 0.23
$\tau\tau\gamma\gamma$ 133 105 93 40 2.50 0.02
@@ IV Sensitivities on the Higgs-gauge boson couplings
where $N_{i}^{B}$ is the number of events of relevant backgrounds and $N_{i}^{TOT}$ is the total number of events including contributions of effective couplings in ith bin of the invariant mass $m_{\ell\ell\gamma\gamma}$ distribution of $\ell^{+}\ell^{-}\gamma\gamma$ system. $\Delta_{i}=\sqrt{\frac{1}{N_{i}^{B}}}$ is the statistical uncertainty in each bin. Fig. 7 shows the invariant mass $m_{\ell\ell\gamma\gamma}$ distribution of $\ell^{+}\ell^{-}\gamma\gamma$ system after all applied cuts for signal and relevant background processes at the muon collider.
@@
Figure 7: The invariant mass $m_{\ell\ell\gamma\gamma}$ distribution of $\ell^{+}\ell^{-}\gamma\gamma$ final state after Cut-5 for signal and relevant background processes at the muon collider.
In Table 3, the 95% C.L. limits of the coefficients $\overline{c}_{\gamma}$, $\overline{c}_{HB}$, $\overline{c}_{HW}$, $\widetilde{c}_{\gamma}$, $\widetilde{c}_{HB}$ and $\widetilde{c}_{HW}$ for center-of-mass energy of 10 TeV with integrated luminosity of 10 ab-1 at muon collider are given and these limits are compared with the limits of experimental studies. At the muon collider, it is noticed the most sensitive limits between all coefficients belong to the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$.
@@
In Section II, we mentioned that the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ make the largest contribution to the total cross-sections as a function of the Wilson coefficients and that these two coefficients are the focus of this study. Table 3 shows that the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ are the most sensitive limits at the muon collider, and however, these two coefficients are even more sensitive than the experimental studies with the best limits. The limits of the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ are discussed in detail in Section V, comparing them with the limits of experimental and phenomenological studies.
@@
Table 3: The 95% C.L. limits for the coefficients $\overline{c}_{\gamma}$, $\overline{c}_{HB}$, $\overline{c}_{HW}$, $\widetilde{c}_{\gamma}$, $\widetilde{c}_{HB}$ and $\widetilde{c}_{HW}$ at muon collider. The limits are compared with the limits of experimental studies.
Coefficients Muon Collider Experimental Limits
$\overline{c}_{\gamma}$ [-0.1001; 0.1216] [-0.00074; 0.00057] Ref. Aad:2016hws
[-0.00011; 0.00011] Ref. ATLAS:2019sdf
$\widetilde{c}_{\gamma}$ [-0.0862; 0.1002] [-0.0018; 0.0018] Ref. Aad:2016hws
[-0.00028; 0.00043] Ref. ATLAS:2019sdf
$\overline{c}_{HB}$ [-0.0052; 0.0040] [-0.086; 0.092] Ref. Aad:2016hws
[-0.057; 0.051] Ref. Aaboud:2018yer
[-0.022; 0.049] Ref. Aaboud:2019ygw
[-0.025; 0.022] Ref. ATLAS:2019sdf
$\widetilde{c}_{HB}$ [-0.2241; 0.2658] [-0.23; 0.23] Ref. Aad:2016hws
[-0.16; 0.16] Ref. Aaboud:2018yer
[-0.065; 0.063] Ref. ATLAS:2019sdf
$\overline{c}_{HW}$ [-0.0034; 0.0023] [-0.086; 0.092] Ref. Aad:2016hws
[-0.057; 0.051] Ref. Aaboud:2018yer
[-0.003; 0.008] Ref. Aaboud:2019ygw
[-0.025; 0.022] Ref. ATLAS:2019sdf
$\widetilde{c}_{HW}$ [-0.1154; 0.0965] [-0.23; 0.23] Ref. Aad:2016hws
[-0.16; 0.16] Ref. Aaboud:2018yer
[-0.065; 0.063] Ref. ATLAS:2019sdf
@@
As discussed above, when five of the six Wilson coefficients can be overly constrained, the one-parameter analysis can be beneficial in sensitivity analysis scenarios. To examine the change in the limits of the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$, which is the focus of this study, we consider the limits where all coefficients except the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ are set to zero. The 95% C.L. contour is obtained while $\chi^{2}$, corresponding to two degrees of freedom, is equal to 5.99 in this analysis. The 95% C.L. contour for the $\overline{c}_{HB}-\overline{c}_{HW}$ plane at the muon collider with the two-parameter analysis is given in Fig. 8.
Figure 8: Two-dimensional 95% C.L. intervals in plane for $\overline{c}_{HB}-\overline{c}_{HW}$ at the muon collider.
@@ V Conclusions
The future muon collider has recently attracted considerable attention due to its multi-TeV energy and clean environment for exploring the Higgs interactions in new physics beyond the SM. At the muon collider with a center-of-mass energy of 10 TeV and an integrated luminosity of 10 ab-1, we examine the dimension-six operators of the Higgs-gauge boson couplings $H\gamma\gamma$, $H\gamma Z$ and $HZZ$ through the process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma$ using the effective Lagrangian approximation in the SILH basis. We analyze the $\ell\ell\gamma\gamma$ production along with the decay of the $Z$-boson by taking into account the effects of a realistic detector with a tuned muon collider detector card. In order to separate the signal events from the backgrounds, we provide the kinematic distributions of the signal and relevant background processes in a cut-based analysis. The 95% C.L. bounds on the Wilson coefficients $\overline{c}_{\gamma}$, $\overline{c}_{HB}$, $\overline{c}_{HW}$, $\widetilde{c}_{\gamma}$, $\widetilde{c}_{HB}$ and $\widetilde{c}_{HW}$ in the SILH basis are obtained using the $\chi^{2}$ test from the invariant mass distribution of the $\ell^{+}\ell^{-}\gamma\gamma$ final state.
@@
In this study, we focus on the sensitivity of the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ at the muon collider. The 95% C.L. bounds on the Wilson coefficients in the $H\rightarrow\gamma\gamma$ decay channel were investigated by the ATLAS collaboration at $\sqrt{s}=13$ TeV with ${\cal L}_{\text{int}}=139$ fb-1, and the experimental limit of the $\overline{c}_{HB}=\overline{c}_{HW}$ coefficients was obtained as [-0.025; 0.022] ATLAS:2019sdf . In the experimental study at $\sqrt{s}=13$ TeV with ${\cal L}_{\text{int}}=79.8$ fb-1, the ATLAS collaboration determined the 95% C.L. bound of the coefficient $\overline{c}_{HW}$ in the $H\rightarrow b\overline{b}$ decay channel as [-0.003; 0.008] Aaboud:2019ygw . Comparing the limits of these experimental studies with our limits at the muon collider, the limit of the coefficient $\overline{c}_{HB}$ is about 5 times more sensitive than in Ref. ATLAS:2019sdf and the limit of the coefficient $\overline{c}_{HW}$ is about 2 times more sensitive than in Ref. Aaboud:2019ygw .
@@
If we consider some phenomenological studies; in Ref. Denizli:2018rca , the 95% C.L. limits of the $\overline{c}_{HB}$ and $\overline{c}_{HW}$ coefficients at the CLIC with $\sqrt{s}=380$ GeV and ${\cal L}_{\text{int}}=500$ fb-1 are [-0.0482; 0.0153] and [-0.00658; 0.00555], respectively, and in Ref. Englert:2016tgb , the 95% C.L. limits of the $\overline{c}_{HB}$ and $\overline{c}_{HW}$ coefficients at the LHC with $\sqrt{s}=14$ TeV and ${\cal L}_{\text{int}}=3000$ fb-1 are [-0.004; 0.004] and [-0.004; 0.004], respectively, in Ref. Shi:2019yhw , the 95% C.L. limits of $\overline{c}_{HB}$ and $\overline{c}_{HW}$ coefficients at the LHC with $\sqrt{s}=13$ TeV and ${\cal L}_{\text{int}}=36.1$ fb-1 are [-0.230; 0.236] and [-0.236; 0.231], respectively, and in Ref. Hesari:2018elr , the 95% C.L. limits of $\overline{c}_{HW}$ coefficient at the FCC-he with $E_{e}=60$ GeV and ${\cal L}_{\text{int}}=10$ ab-1 are found to be [-0.017; 0.019]. According to these phenomenological studies, the limit of coefficient $\overline{c}_{HB}$ is about 7 times and 50 times more sensitive than in Ref. Denizli:2018rca and in Ref. Shi:2019yhw , respectively, and almost the same order as in Ref. Englert:2016tgb , and the limit of coefficient $\overline{c}_{HW}$ is about 2 times, 82 times and 6 times more sensitive than in Ref. Denizli:2018rca , in Ref. Shi:2019yhw and in Ref. Hesari:2018elr , respectively, and almost the same order as in Ref. Englert:2016tgb .

### v2

@@ III Analysis of Signal and Background Events
$\overline{c}_{\gamma}=0.5$ 123 0.043 120 0.472 111 0.539 77 0.696 66 1.183 23 17.35
$\widetilde{c}_{\gamma}=0.5$ 174 0.061 170 0.669 152 0.737 106 0.959 92 1.649 33 24.91
$\overline{c}_{HB}=0.5$ 10867 3.813 10623 41.82 8900 43.20 8493 76.83 7530 134.9 6461 4876
$\widetilde{c}_{HB}=0.5$ 98 0.034 96 0.377 84 0.407 52 0.470 44 0.788 4.37 3.298
$\overline{c}_{HW}=0.5$ 25086 8.802 24722 97.33 21780 105.7 20847 188.6 18695 335.1 15605 11777
$\widetilde{c}_{HW}=0.5$ 111 0.039 109 0.429 97 0.471 66 0.597 56 1.003 18 13.58
Backgrounds
$Z\gamma\gamma$ 70 69 60 35 30 0.27
$\tau\tau\gamma\gamma$ 133 119 86 36 12 0.06
ZZZZ 59 26 25 20 4.87 0.14
$t\bar{t}\gamma\gamma$ 23 17 14 5.54 1.20 0.005
ZjjZjj 2565 23 21 14 7.71 0.85
@@ IV Sensitivities on the Higgs-gauge boson couplings
where $N_{i}^{B}$ is the number of events of relevant backgrounds and $N_{i}^{TOT}$, as a function of Wilson coefficients, is the total number of events including contributions of effective couplings in ith bin of the invariant mass $m_{\ell\ell\gamma\gamma}$ distribution of $\ell^{+}\ell^{-}\gamma\gamma$ system. $\Delta_{i}=\sqrt{\delta_{sys}^{2}+\delta_{st}^{2}}$ includes systematic uncertainty $(\delta_{sys})$ and statistical uncertainty in each bin $(\delta_{st}=1/\sqrt{N_{i}^{B}})$ Spor:2022oer . Fig. 7(a) and Fig. 7(b) show the invariant mass $m_{\ell\ell\gamma\gamma}$ distribution of $\ell^{+}\ell^{-}\gamma\gamma$ system after Cut-0 (left) and after all applied cuts (right) for the signal and relevant background processes at the muon collider.
@@
Figure 7: The invariant mass $m_{\ell\ell\gamma\gamma}$ distribution of $\ell^{+}\ell^{-}\gamma\gamma$ final state after Cut-0 (left) and after Cut-5 (right) for signal and relevant background processes at the muon collider.
Although the center-of-mass energy is 10 TeV, the process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma\rightarrow\ell\ell\gamma\gamma$ is not a direct 10 TeV resonance production. The $Z$-boson and photons produced in this process do not carry all the center-of-mass energy; most of the event is carried by particles and radiation traveling in very forward directions. In high-energy collisions, some particles and radiation are scattered along the beamline with small angles and cannot be detected by forward detectors if they are outside the acceptance range, thus increasing energy losses. In particular, a significant amount of energy may not be captured by the detector due to the initial-state radiation (ISR) and final-state radiation (FSR). Although muons radiate less than electrons, these effects cannot be ignored at very high energies, such as 10 TeV. In short, the reason why the observed $m_{\ell\ell\gamma\gamma}$ distribution does not peak at 10 TeV is that due to the nature of the process, the total energy of the observed final-state particles is much lower than the center-of-mass energy in events and the unobserved radiation plays an important role.
The median expected significance $Z$ for the signal discovery is computed as Cowan:2011wer ; Kumar:2015tgv ; Senol:2024umn :
$Z=\sqrt{2[(S+B)\text{ln}(\frac{(S+B)(1+\delta^{2}B)}{B+\delta^{2}B(S+B)})-\frac{1}{\delta^{2}}\text{ln}(1+\frac{\delta^{2}S}{1+\delta^{2}B})]}$ (15)
where $\delta$ is the systematic uncertainty. $S$ and $B$ are the numbers of events obtained by the invariant mass $m_{\ell\ell\gamma\gamma}$ distribution of $\ell^{+}\ell^{-}\gamma\gamma$ system for the signal and the sum of the relevant backgrounds, respectively. In the case where the background expectation is known exactly, with the limit of $\delta\rightarrow 0$, this expression simplifies to:
$Z=\sqrt{2[(S+B)\text{ln}(1+S/B)-S]}$ (16)
@@
In particle physics, the $5\sigma$” standard is used to define discovery sensitivity, that is, the statistical significance must be $Z>5$; in the case of the $3\sigma$” standard ($Z>3$), it is claimed that the observed signal has strong evidence. In Table 3, the 95% C.L., $3\sigma$ and $5\sigma$ limits of the coefficients $\overline{c}_{\gamma}$, $\overline{c}_{HB}$, $\overline{c}_{HW}$, $\widetilde{c}_{\gamma}$, $\widetilde{c}_{HB}$ and $\widetilde{c}_{HW}$ for the center-of-mass energy of 10 TeV with integrated luminosity of 10 ab-1 at muon collider are given without and with systematic uncertainty $\delta=15\%$. The most sensitive limits between all coefficients belong to the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$. In Section II, we mentioned that the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ make the largest contribution to the total cross-sections as a function of the Wilson coefficients and that these two coefficients are the focus of this study. Table 3 shows that the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ are the most sensitive limits at the muon collider, and however, the 95% C.L. limits of these two coefficients are even more sensitive than the experimental studies with the best limits. The limits of the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ are discussed in detail in Section V, comparing them with the limits of experimental and phenomenological studies.
@@
Table 3: The limits at 95% C.L., $3\sigma$ and $5\sigma$ for the coefficients $\overline{c}_{\gamma}$, $\overline{c}_{HB}$, $\overline{c}_{HW}$, $\widetilde{c}_{\gamma}$, $\widetilde{c}_{HB}$ and $\widetilde{c}_{HW}$ without and with systematic uncertainty $\delta=15\%$ at muon collider.
Coefficients $\delta_{sys}$ $\chi^{2}$ at 95% C.L. $Z$ at $3\sigma$ $Z$ at $5\sigma$
$\overline{c}_{\gamma}$ 0 [-0.1103; 0.1392] [-0.2243; 0.2236] [-0.3052; 0.3027]
15% [-0.2259; 0.2151] [-0.4465; 0.4478] [-0.8791; 0.8801]
$\widetilde{c}_{\gamma}$ 0 [-0.1021; 0.1080] [-0.2029; 0.2052] [-0.2773; 0.2724]
15% [-0.2016; 0.2026] [-0.3742; 0.3819] [-0.6245; 0.7989]
$\overline{c}_{HB}$ 0 [-0.0058; 0.0057] [-0.0116; 0.0115] [-0.0162; 0.0161]
15% [-0.0111; 0.0118] [-0.0255; 0.0254] [-0.0365; 0.0366]
$\widetilde{c}_{HB}$ 0 [-0.2588; 0.3019] [-0.5125; 0.4757] [-0.6661; 0.5970]
15% [-0.5505; 0.4839] [-1.0174; 0.9116] [-1.8689; 1.8369]
$\overline{c}_{HW}$ 0 [-0.0047; 0.0035] [-0.0075; 0.0074] [-0.0114; 0.0113]
15% [-0.0071; 0.0075] [-0.0178; 0.0179] [-0.0251; 0.0252]
$\widetilde{c}_{HW}$ 0 [-0.1342; 0.1212] [-0.2574; 0.2658] [-0.3420; 0.3741]
15% [-0.2563; 0.2615] [-0.5186; 0.5250] [-0.9334; 0.9331]
@@
As discussed above, when five of the six Wilson coefficients can be overly constrained, the one-parameter analysis can be beneficial in sensitivity analysis scenarios. To examine the change in the limits of the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$, which is the focus of this study, we consider the limits where all coefficients except the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ are set to zero. The 95% C.L. contour is obtained while $\chi^{2}$, corresponding to two degrees of freedom, is equal to 5.99 in this analysis. The 95% C.L. contour for the $\overline{c}_{HB}-\overline{c}_{HW}$ plane at the muon collider using the two-parameter analysis is given without and with systematic uncertainty $\delta=15\%$ in Fig. 8(a).
The statistical significance ($Z$) as a function of the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ of the Higgs-gauge boson couplings without and with systematic uncertainty $\delta=15\%$ is shown in Fig. 8(b). From this figure, the bounds on the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ correspond to the intersection of the curves with the horizontal blue and green lines representing the $3\sigma$ and $5\sigma$ levels, respectively.
Figure 8: Two-dimensional 95% C.L. intervals in plane for $\overline{c}_{HB}-\overline{c}_{HW}$ (left) and the statistical significance ($Z$) as a function of the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ (right).
The validity of the EFT can be tested by examining the relationship between the new physics scale and the Wilson coefficients of dimension-six operators as $\bar{c}\sim g^{2}_{*}\nu^{2}/\Lambda_{s}^{2}$ Ellis:2015uja ; Contino:2013ert ; Karadeniz:2020wlo . Here, $\nu$ is the vacuum expectation value of the Higgs field and $g_{*}$ is the coupling constant of the heavy degrees of freedom with the SM particles. The upper bound on the new physics scale with $\nu=246$ GeV and $g_{*}=4\pi$ for strongly-coupled new physics are 40.9 TeV for $\overline{c}_{HB}$:[-0.0058; 0.0057] and 45.1 TeV for $\overline{c}_{HW}$:[-0.0047; 0.0035]. In this study, since the energy scale depends on the specific process and the observable, $E\sim m_{\ell\ell\gamma\gamma}$, these upper bounds are within the validity range of the EFT. An uncertainty arises from the truncation of the EFT expansion because it neglects the contribution from dimension-eight, which may be comparable to or larger than the contribution from dimension-six. However, including the full set of dimension-eight operators is often quite complex; therefore, the SMEFT with only dimension-six effective operators is typically used, but attention should be paid to the potential failure of this truncation and the possible relative contribution of some dimension-eight terms. In this study, the potential truncation error is considered within the systematic uncertainty.
@@ V Conclusions
The future muon collider has recently attracted considerable attention due to its multi-TeV energy and clean environment for exploring the Higgs interactions in new physics beyond the SM. At the muon collider with a center-of-mass energy of 10 TeV and an integrated luminosity of 10 ab-1, we examine the dimension-six operators of the Higgs-gauge boson couplings $H\gamma\gamma$, $H\gamma Z$ and $HZZ$ through the process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma$ using the effective Lagrangian approximation in the SILH basis. We analyze the $\ell\ell\gamma\gamma$ production along with the decay of the $Z$-boson by taking into account the effects of a realistic detector with a tuned muon collider detector. In order to separate the signal events from the backgrounds, we provide the kinematic distributions of the signal and relevant background processes in a cut-based analysis. The 95% C.L., $3\sigma$ and $5\sigma$ bounds on the Wilson coefficients $\overline{c}_{\gamma}$, $\overline{c}_{HB}$, $\overline{c}_{HW}$, $\widetilde{c}_{\gamma}$, $\widetilde{c}_{HB}$ and $\widetilde{c}_{HW}$ in the SILH basis are obtained without and with systematic uncertainty of 15% using the $\chi^{2}$ test and statistical significance test from the invariant mass distribution of the $\ell^{+}\ell^{-}\gamma\gamma$ final state.
@@
In this study, we focus on the sensitivity of the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ at the muon collider. In the experimental study at $\sqrt{s}=13$ TeV with ${\cal L}_{\text{int}}=79.8$ fb-1, the ATLAS collaboration determined the 95% C.L. bound of the coefficient $\overline{c}_{HB}$ and $\overline{c}_{HW}$ in the $H\rightarrow b\overline{b}$ decay channel as [-0.022; 0.049] and [-0.003; 0.008], respectively Aaboud:2019ygw . The 95% C.L. bounds on the Wilson coefficients in the $H\rightarrow\gamma\gamma$ decay channel were investigated by the ATLAS collaboration at $\sqrt{s}=13$ TeV with ${\cal L}_{\text{int}}=139$ fb-1, and the experimental limit of the $\overline{c}_{HB}=\overline{c}_{HW}$ coefficients was obtained as [-0.025; 0.022] ATLAS:2019sdf . Among the limits of these two experimental studies, the most sensitive limit of the coefficient $\overline{c}_{HB}$ is in Ref. ATLAS:2019sdf and the most sensitive limit of the coefficient $\overline{c}_{HW}$ is in Ref. Aaboud:2019ygw . Comparing the limits of these experimental studies with our limits at 95% C.L., the limit of the coefficient $\overline{c}_{HB}$ is about 4.1 times more sensitive than in Ref. ATLAS:2019sdf . Our limit at 95% C.L. of the coefficient $\overline{c}_{HW}$ is about 1.3 times more sensitive than in Ref. Aaboud:2019ygw .
@@
If we consider some phenomenological studies; in Ref. Denizli:2018rca , the 95% C.L. limits of the $\overline{c}_{HB}$ and $\overline{c}_{HW}$ coefficients at the CLIC with $\sqrt{s}=380$ GeV and ${\cal L}_{\text{int}}=500$ fb-1 are [-0.0482; 0.0153] and [-0.00658; 0.00555], respectively, and in Ref. Englert:2016tgb , the 95% C.L. limits of the $\overline{c}_{HB}$ and $\overline{c}_{HW}$ coefficients at the LHC with $\sqrt{s}=14$ TeV and ${\cal L}_{\text{int}}=3000$ fb-1 are [-0.004; 0.004] and [-0.004; 0.004], respectively, in Ref. Shi:2019yhw , the 95% C.L. limits of $\overline{c}_{HB}$ and $\overline{c}_{HW}$ coefficients at the LHC with $\sqrt{s}=13$ TeV and ${\cal L}_{\text{int}}=36.1$ fb-1 are [-0.230; 0.236] and [-0.236; 0.231], respectively, and in Ref. Hesari:2018elr , the 95% C.L. limits of $\overline{c}_{HW}$ coefficient at the FCC-he with $E_{e}=60$ GeV and ${\cal L}_{\text{int}}=3$ ab-1 are found to be [-0.0032; 0.0035]. According to these phenomenological studies, the limits at 95% C.L. of the coefficient $\overline{c}_{HB}$ are more sensitive than in Ref. Denizli:2018rca and in Ref. Shi:2019yhw and more stringent than in Ref. Englert:2016tgb . The limits at 95% C.L. of the coefficient $\overline{c}_{HW}$ are more sensitive than in Ref. Denizli:2018rca and in Ref. Shi:2019yhw and more stringent than in Ref. Englert:2016tgb and Ref. Hesari:2018elr . The comparison of 95% C.L. limits of the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ in this study with the limits of present experimental and various phenomenological studies is summarized in Table 4.
Table 4: Comparison of the 95% C.L. limits of $\overline{c}_{HB}$ and $\overline{c}_{HW}$ coefficients in this study with present experimental and phenomenological studies.
Coefficients $\overline{c}_{HB}$ $\overline{c}_{HW}$
This Study $\chi^{2}\rightarrow 95\%$ C.L. [-0.0058; 0.0057] [-0.0047; 0.0035]
Experimental Ref. Aaboud:2019ygw [-0.022; 0.049] [-0.003; 0.008]
Limits (ATLAS) Ref. ATLAS:2019sdf [-0.025; 0.022] [-0.025; 0.022]
Ref. Denizli:2018rca [-0.0482; 0.0153] [-0.00658; 0.00555]
Phenomenological Ref. Englert:2016tgb [-0.004; 0.004] [-0.004; 0.004]
Limits Ref. Shi:2019yhw [-0.230; 0.236] [-0.236; 0.231]
Ref. Hesari:2018elr [-0.0032; 0.0035]
@@
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.